# Natural Language Processing in Patents: A Survey

### v2

@@ 4.2 Curated Data Collections
Datasets. Patent offices provide large-scale raw data in the patent domain. Developers and researchers rely on well-curated datasets for development and research. We summarize representative curated patent datasets in Table 2. We aim to reduce the time and effort for data searching in the patent domain by presenting these ready-to-use datasets, especially the publicly available ones.
@@
The Japanese National Institute of Informatics Testbeds and Community for Information access Research (NTCIR) [^10] NTCIR: https://research.nii.ac.jp/ntcir/index-en.html provides datasets and organizes shared tasks to facilitate research in information retrieval, natural language processing, and related areas. Patent-related tasks at the NTCIR range from patent retrieval and classification to text mining and machine translation (Lupu et al., 2017).
TREC-CHEM [^11] TREC-CHEM: https://trec.nist.gov/data/chem-ir.html is a part of the Text REtrieval Conference (TREC) series, specialized on chemical information retrieval. It contains patent datasets that are rich in chemical information. For example, the dataset from TREC-CHEM 2009 contains 2.6 million patent files registered at the European Patent Office, United States Patent and Trademark Office, and World Intellectual Property Organization (Lupu et al., 2009).
@@ 5.1.4 Methodologies for Automated Subject Classification
Figure 6: Three mainstream methods for patent classification
@@ 5.9.1 Task Definition of Innovation Recommendation
Innovation recommendation refers to the process of suggesting new ideas or methods in a technology context. The automation of this task can enhance competitiveness, efficiency, or user experience. Innovation recommendations should help companies to identify novel research and development opportunities.181818 Strictly speaking, the use of AI to generate inventions may raise legal issues, notably around inventorship and ownership. The use of AI for writing a patent could do so if the AI starts to add information or hallucinate so that it could be considered inventive, not considering copyright and authorship rights. Traditionally, patents are granted to human inventors, i.e., natural persons. The US Patent and Trademark Office, at present, does not consider AI-generated inventions as patentable, which limits how technology gap detection and idea generation can be used to create patentable inventions. Practically, natural co-inventors might suppress the co-inventorship of an AI until technical means are available to detect AI contributions. No precedence seems available if that would collide the inventors’ oath filed with the application or would set an inventor into bad faith. These issues highlight the need for evolving legal standards to accommodate the rapid advancements of AI technology in the patent system.
Figure 13: Methods of innovation recommendation task
@@ 6 Patent Generation Tasks
Figure 14: Illustration of text generation tasks
@@ 6.1.1 Task Definition of Summarization
Patent summarization aims to create concise and informative summaries of patent documents. Given the complex and technical nature of patents, summarization helps in extracting the most important information and paraphrasing in more accessible language depending on the audience. The target audience may be as diverse as patent examiners, researchers, development engineers, and legal professionals. This process involves understanding and condensing the key aspects of a patent, such as claims, background, and detailed descriptions of the invention.
@@ 6.1.2 Methodologies for Summarization
LLMs have achieved satisfactory performance on most summarization tasks, even surpassing the benchmark of reference summaries (Pu et al., 2023). However, how these LLMs, such as GPT-4, perform on patent summarization is still worth investigating. The technical content and legal information in patents may pose some challenges to LLMs compared to normal summarization.
@@ 6.2.2 Methodologies for Translation
Fixing Specific Errors. Since accurate language is crucial in patent texts while language models are even designed to vary based on distance metrics of word meanings, for instance, research has studied pre- and post-processing methods for fixing common problems. As an example for terminology errors, Ying et al. (2021) identified eight types of terminology errors in patent translation from English into Chinese and suggested solving these errors by pre-editing the source texts. In addition, Larroyed (2023) compared ChatGPT (Ray, 2023) and the Patent Translate system from the European Patent Office on patent translation task. In their analysis, Patent Translate could use its training on patent corpora and specialization in patent translation to outperform ChatGPT in accuracy. ChatGPT, however, performed better on language structure and overall textual coherence. In linguistics, these two qualities might refer to paradigmatic versus syntagmatic properties. Such results suggest that a combination of both advantages may open new opportunities. One part of the solution may be on the model side, where transformers or a combination of recurrent neural networks (RNNs) and transformers may improve structural, i.e., syntagmatic shortcomings. Further deriving a domain-specific model through training and/or fine-tuning to patent corpora promises to enhance both syntagmatic and paradigmatic performance through the incorporation of more patent-typical language structures as well as sharper control over terminology-related issues.
@@ 6.3.2 Methodologies for Simplification
In contrast to well-established tasks of patent summarization and translation, patent simplification is less investigated and worth more attention. LLMs appear an ideal tool for this task, despite the challenges of processing complex technical documents with legal requirements.
@@ 6.4 Patent Writing
Table 5: Previous studies for patent writing task.
Paper Task Method Data
Lee and Hsiang (2020b) Claim generation Fine-tunning GPT-2 555,890 patent claims from granted U.S. utility patents in 2013 (Claims are the first and independent claims)
Lee (2020) Title, abstract, and claim generation Fine-tuning transformer-based models 404,975,822 utility patent from 1976 to 2017-08 in Google Patents Public Datasets
Jiang et al. (2024) Description-based claim generation Fine-tuning Llama-3-8B, GPT-4 (zero-shot) 9,500 pairs of patent descriptions and claims
Christofidellis et al. (2022) Claim generation, text infilling Fine-tuning GPT-2 with Multi-task Learning 11,600,000 patents published between 1998 and 2020
@@ 6.4.2 Methodologies for Patent Writing
A few studies have investigated patent writing for different sections based on fine-tuning LLMs. Lee and Hsiang (2020a) provided an initial proof-of-concept study for patent claim generation. The authors fine-tuned the GPT-2 model to generate patent claims, but the quality of generated claims was not measured, which reduced the actual value of the work. In further research, Lee (2020) trained the GPT-2 model to map one section to another, such as generating abstracts based on title and generating claims from abstracts. Nonetheless, patent professionals may raise concerns about the validity of the tasks. Patent titles only include a few words and lack specificity, which makes title-based generation unfounded. In contrast, abstract-based claim generation sometimes makes sense because the abstract in some patents is just a paraphrased version of the first independent claim (without the legal phraseology). Hence, this task is straightforward by extracting and revising the abstract. However, abstracts are typically crafted to be general and reveal minimal details of the invention, adhering to requirements of patent offices (European Patent Office, 2000; World Intellectual Property Organization, 2022). Consequently, it is almost impossible to generate detailed invention features required by claims based solely on general and vain abstracts. Therefore, both title-based and abstract-based generation tasks are not well-conditioned. Alternatively, patent descriptions include all details and specific embodiments of the invention, which makes description-based patent generation notably promising for future research. Jiang et al. (2024) constructed a dataset with 9,500 examples and evaluated the performance of various LLMs in patent claim generation. The results showed that description-based claim generation outperformed previous research relying on abstracts. Moreover, fine-tuning can enhance the claims’ completeness, conceptual clarity, and logical linkage. Despite promising capabilities, comprehensive revisions are still necessary for LLM-generated claims to pass rigorous patent scrutiny.
@@
In addition, Christofidellis et al. (2022) proposed a prompt-based generative transformer for the patent domain, which used GPT-2 as backbone and trained with multi-task learning (Maurer et al., 2016) on part-of-patent generation, text infilling, and patent coherence evaluation tasks. Specifically, the training processes involved two text generation tasks, generating patent titles based on abstracts and suggesting words for masked tokens in the given abstracts. The results showed that the performance was better than other baselines, such as BERT and GPT-2 on a single task. This paper indicates that multi-task learning may be a promising method for patent text generation in the future.
@@ 7.1 Data and Benchmarks
Additionally, benchmarks consist of labeled datasets and established metrics for performance evaluation, such as accuracy and precision. Benchmarks provide a standard and fair comparison platform for different methods. In the patent domain, many studies evaluated their approaches based on closed-source datasets, and there is a lack of effective evaluation metrics for patent text generation. Well-defined benchmarks can ensure the results are reproducible and comparable.
@@ 7.2 Use of Large Language Models
Since texts are the key ingredient to patent documents, promising methods to analyze and generate patent texts are worth investigating. Large language models (LLMs) have dramatically changed the field and achieved remarkable performance in a wide range of tasks in the general domain (Min et al., 2023). However, few works have evaluated and investigated the usage of LLMs on patent-related tasks, leaving a large research gap. LLMs could identify more intricate and nuanced patterns and relationships within patent texts, facilitating the effectiveness in patent analysis tasks, such as patent classification and novelty prediction. Furthermore, LLMs could assist in generating and refining patent texts based on a comprehensive understanding of existing patent literature. The integration of LLMs into patent tasks cannot only enhance the efficiency and effectiveness of patent analysis but also improve the quality of the patent drafting process. As these models continue to evolve, LLMs appear significantly promising in the patent domain.
@@ 7.5 Multimodal Methods
Apart from texts, patent documents themselves and their context include other useful information, such as drawings and citations. Multimodal models can combine multiple data types to improve the performance of patent analysis. Currently, only a few researchers have used this approach, for example in patent classification (Jiang et al., 2022a). Ideally, researchers can adopt multimodal methods in other patent tasks as well. For example, multimodal models can assist in patent drafting from design figures, ensuring the textual descriptions and figures are aligned. Furthermore, multimodal methods can create more informative and intuitive visualizations of patent data, providing a clearer understanding of the patent landscape. A multimodal approach appears promising and deserves more attention.
@@ 8 Conclusion
NLP techniques play a prominent role in the automation of patent processing tasks. Text embeddings can extract word and semantic information for patent analysis. Sentence/paragraph embeddings are preferable to word embeddings because they can capture more nuanced contextual information from entire sentences or paragraphs. From the model perspective, most traditional machine-learning models, such as feature-based neural networks, support vector machines, or random forests, have already been outdated. Deep learning models, particularly convolutional and recurrent neural networks, are still used in some circumstances, for example when the computing power is limited. However, larger transformer-based models, such as the BERT and the GPT series, have almost revolutionized the field despite their excessive need for computational resources. Furthermore, scaling the model size to LLMs has led to an explosion in performance on numerous general tasks. However, most LLMs adapted and tested in the patent domain are still on the level of GPT-2 (Radford et al., 2019) with approximately 1.5 billion parameters.
@@
Research has prominently focused on short text parts of patents, such as titles and abstracts, e.g., for patent analysis or generation tasks. However, these texts are the least specific texts with little information about the actual invention. Since they are highly generic and take little time during drafting, the automation of these short texts is not necessarily helpful. Among the shorter parts of patents, the claims would be an exception, defining the legal scope of patent protection. The task of claim generation is worth investigating (Jiang et al., 2024).
@@
The most studied and developed topics are automated patent classification and retrieval, which are routine tasks in patent offices. Transformer-based language models, such as BERT, have shown better effectiveness in leveraging text information compared to traditional machine learning models and deep word embeddings. A more promising direction is to hybridise methods so that they can use multiple sources from patents, including texts, images, and metadata. Information extraction serves as a primer to support various patent-related applications. State-of-the-art language models can help in precise extraction to improve the performance in further applications. Furthermore, we found that the conception of novelty used in research can substantially deviate from its strict definition in patent law. Accordingly, a share of the studies may aim at predicting the patent novelty but actually generate some form of metric for oddness, potential creativity, or perceived originality. Such cases indicate the often underestimated difficulties when working with patents. Interestingly, there are practically no capable foundation models that could deal with the particularities of the patent domain and simplify the derivation of usable tools for specific tasks. Granting prediction, litigation prediction, and patent valuation are comparably less investigated. The key acts in patent examination are almost ignored so far, including the automated formal novelty and inventiveness detection for higher quality and objectivity, deriving of arguments, and support in formulating motions.
@@
Latest language models play a role in generative tasks around patents and reflect the most recent trend towards generative artificial intelligence technology. Apart from the patent analysis tasks, the language particularities of patents also complicate generative tasks. Summarization tasks and automated translation tasks are well established and available online. Particularly, automated translation has been studied and in production at patent offices for more than a decade. Less conspicuous yet socially beneficial tasks, such as patent simplification, receive scant attention in research. This oversight is primarily attributed to the absence of appropriate datasets. Although patent drafting, especially the sections on descriptions and claims, is an obvious task, the available research is limited and generally unsatisfactory.

### v3

@@ 4.2 Curated Data Collections
Datasets. Patent offices provide large-scale raw data in the patent domain. Developers and researchers rely on well-curated datasets for development and research. We summarize representative publicly available curated patent datasets in Table 2. We aim to reduce the time and effort for data searching in the patent domain by presenting these ready-to-use datasets.
@@
The Japanese National Institute of Informatics Testbeds and Community for Information access Research (NTCIR) [^11] NTCIR: https://research.nii.ac.jp/ntcir/index-en.html provides datasets and organizes shared tasks to facilitate research in information retrieval, natural language processing, and related areas. Patent-related tasks at the NTCIR range from patent retrieval and classification to text mining and machine translation (Utiyama and Isahara, 2007; Lupu et al., 2017).
TREC-CHEM [^12] TREC-CHEM: https://trec.nist.gov/data/chem-ir.html is a part of the Text REtrieval Conference (TREC) series, specializing in chemical information retrieval. It contains patent datasets that are rich in chemical information. For example, the dataset from TREC-CHEM 2009 contains 2.6 million patent files registered at the European Patent Office, United States Patent and Trademark Office, and World Intellectual Property Organization (Lupu et al., 2009).
@@ 5.1.2 Classification Scheme
Figure 6: Three mainstream methods for patent classification
@@ 5.9.1 Task Definition of Innovation Recommendation
Innovation recommendation refers to the process of suggesting new ideas or methods in a technology context. The automation of this task can enhance competitiveness, efficiency, or user experience. Innovation recommendations should help companies to identify novel research and development opportunities.191919 Strictly speaking, the use of AI to generate inventions may raise legal issues, notably around inventorship and ownership. The use of AI for writing a patent could do so if the AI starts to add information or hallucinate so that it could be considered inventive, not considering copyright and authorship rights. Traditionally, patents are granted to human inventors, i.e., natural persons. The US Patent and Trademark Office, at present, does not consider AI-generated inventions as patentable, which limits how technology gap detection and idea generation can be used to create patentable inventions. Practically, natural co-inventors might suppress the co-inventorship of an AI until technical means are available to detect AI contributions. No precedence seems available if that would collide the inventors’ oath filed with the application or would set an inventor into bad faith. These issues highlight the need for evolving legal standards to accommodate the rapid advancements of AI technology in the patent system.
Figure 14: Illustration of text generation tasks
@@ 6 Patent Generation Tasks
Model Architecture R-1 R-2 R-L
Seq2Seq (Sharma et al., 2019) LSTM 28.74 7.87 24.66
SentRewriting (Sharma et al., 2019) LSTM 37.12 11.87 32.45
PEGASUS (Zhang et al., 2020) Transformer 43.55 20.43 31.80
BigBird-PEGASUS (Zaheer et al., 2020) Transformer 60.64 42.46 50.01
LongT5 (Guo et al., 2022) Transformer 76.87 66.06 70.76
Table 5: Comparison of representative models on the BigPatent benchmark (Sharma et al., 2019) for patent summarization as measured by ROUGE score (Lin, 2004)
@@ 6.1.1 Task Definition of Summarization
Patent summarization aims to create concise and informative summaries of patent documents. Given the complex and technical nature of patents, summarization helps in extracting the most important information and paraphrasing in more accessible language depending on the audience. The target audience may be as diverse as patent examiners, researchers, development engineers, and legal professionals. Additionally, summarization can extract the gist of patent subgroups within classification systems, which is essential for organizing and managing large patent datasets (Souza et al., 2021). The summarization process involves understanding and condensing the key aspects of a patent, such as claims, background, and detailed descriptions of the invention.
@@ 6.1.2 Methodologies for Summarization
Table 5 compares representative models on the BigPatent benchmark (Sharma et al., 2019) for patent summarization. Transformer-based models outperform previous LSTM-based methods. In addition, models with long input lengths, such as LongT5 (Guo et al., 2022), significantly increase the performance, because patent descriptions usually contain more than 10,000 tokens (Suzgun et al., 2023).
LLMs have achieved satisfactory performance on most summarization tasks in the general domain, even surpassing the benchmark of reference summaries (Pu et al., 2023). The state-of-the-art method for text summarization is to use LLMs for abstractive summarization (Rao et al., 2024). However, how these LLMs, such as GPT-4, perform on patent summarization is still worth investigating. The technical content and legal information in patents may pose some challenges to LLMs compared to normal summarization.
@@ 6.2.2 Methodologies for Translation
Fixing Specific Errors. Since accurate language is crucial in patent texts while language models are even designed to vary based on distance metrics of word meanings, for instance, research has studied pre- and post-processing methods for fixing common problems. As an example for terminology errors, Ying et al. (2021) identified eight types of terminology errors in patent translation from English into Chinese and suggested solving these errors by pre-editing the source texts. In addition, Larroyed (2023) compared ChatGPT (Ray, 2023) and the Patent Translate system of the European Patent Office on patent translation task. In their analysis, Patent Translate could exploit its training on patent corpora and specialization in patent translation to outperform ChatGPT in accuracy. ChatGPT, however, performed better on language structure and overall textual coherence. In linguistics, these two qualities might refer to paradigmatic versus syntagmatic properties. Such results suggest that a combination of both advantages may open new opportunities. One part of the solution may be on the model side, where transformers or a combination of recurrent neural networks (RNNs) and transformers may improve structural, i.e., syntagmatic shortcomings. Further deriving a domain-specific model through training and/or fine-tuning to patent corpora promises to enhance both syntagmatic and paradigmatic performance through the incorporation of more patent-typical language structures as well as sharper control over terminology-related issues.
While translation systems, such as Patent Translate, offer valuable tools for understanding patent documents in various languages, it is important to recognize that machine translations may not always achieve the precision required for legal or formal purposes. Therefore, for critical applications such as patent filings or legal proceedings, consultation with professional human translators is advisable to ensure the highest level of accuracy. From the above studies, we can conclude that training specific LLMs on extensive patent translation datasets is highly promising to further improve translation accuracy and language precision.
@@ 6.3.2 Methodologies for Simplification
In contrast to well-established tasks of patent summarization and translation, patent simplification is less investigated and worth more attention. Particularly, established benchmarks to compare different models and approaches are important to accelerate the development of this topic. LLMs appear an ideal tool for this task, despite the challenges of processing complex technical documents with legal requirements.
Task Method Data
Claim generation (Lee and Hsiang, 2020b) Fine-tunning GPT-2 555,890 patent claims from granted U.S. utility patents in 2013 (Claims are the first and independent claims)
Title, abstract, and claim generation (Lee, 2020) Fine-tuning transformer-based models 404,975,822 utility patent from 1976 to 2017-08 in Google Patents Public Datasets
Description-based claim generation (Jiang et al., 2024c) Fine-tuning Llama-3-8B, GPT-4 (zero-shot) 9,500 pairs of patent descriptions and claims
Claim generation, text infilling (Christofidellis et al., 2022) Fine-tuning GPT-2 with Multi-task Learning 11,600,000 patents published between 1998 and 2020
Claim revision (Jiang et al., 2024b) Fine-tuning Llama-3.1-8B, GPT-4 (zero-shot) 22.6K pairs of application-level and granted versions of patent claims
Specification generation (Wang et al., 2024a) Fine-tuning GPT-J and T5 13,725 patents from USPTO in CPC code of G06N
Paper to patent generation (Knappich et al., 2024) Fine-tuning Llama-3-8B 1,800 pairs of academic papers and corresponding patent documents (outlines)
Table 6: Previous studies for patent writing task.
@@ 6.4.2 Methodologies for Patent Writing
A few studies have investigated patent writing for different sections based on fine-tuning LLMs. Lee and Hsiang (2020a) provided an initial proof-of-concept study for patent claim generation. The authors fine-tuned the GPT-2 model to generate patent claims, but the quality of generated claims was not measured, which reduced the actual value of the work. In further research, Lee (2020) trained the GPT-2 model to map one section to another, such as generating abstracts based on title and generating claims from abstracts. Nonetheless, patent professionals may raise concerns about the validity of the tasks. Patent titles only include a few words and lack specificity, which makes title-based generation unfounded. In contrast, abstract-based claim generation sometimes makes sense because the abstract in some patents is just a paraphrased version of the first independent claim (without the legal phraseology). Hence, this task is straightforward by extracting and revising the abstract. However, abstracts are typically crafted to be general and reveal minimal details of the invention, adhering to requirements of patent offices (European Patent Office, 2000; World Intellectual Property Organization, 2022). Consequently, it is almost impossible to generate detailed invention features required by claims based solely on general and vain abstracts. Therefore, both title-based and abstract-based generation tasks are not well-conditioned. Alternatively, patent descriptions include all details and specific embodiments of the invention, which makes description-based patent generation notably promising for future research. Recent research constructed a dataset with 9,500 examples and evaluated the performance of various LLMs in patent claim generation. The results demonstrated that description-based claim generation outperformed previous research, which relied on abstracts (Jiang et al., 2024c). Moreover, fine-tuning can enhance the claims’ completeness, conceptual clarity, and logical linkage. GPT-4 achieved the best performance among all tested LLMs and uniquely grouped alternative embodiments within dependent claims logically. Despite promising capabilities, comprehensive revisions are still necessary for LLM-generated claims to pass rigorous patent scrutiny. Follow-up work further explored the ability of LLMs to revise patent claims. The findings indicate that LLMs often introduce ineffective edits that deviate from the intended revisions, whereas fine-tuning can improve performance (Jiang et al., 2024b). Conversely, Wang et al. (2024a) appear to be the first to investigate the generation of patent specifications from claim and drawing texts. They fine-tuned GPT-J (Wang, 2021) and T-5 (Raffel et al., 2020) and demonstrated their ability to produce human-like patent specifications that adhere to a legal writing style. However, their approach still relies on certain simplifying assumptions that deviate from real-world patent drafting practices, such as the assumption that each specification paragraph corresponds to only one claim feature and a single drawing. While their work lays a solid foundation for specification generation, significant challenges remain before achieving practical applicability.
@@
In addition, Christofidellis et al. (2022) proposed a prompt-based generative transformer for the patent domain, which used GPT-2 as backbone and trained with multi-task learning (Maurer et al., 2016) on part-of-patent generation, text infilling, and patent coherence evaluation tasks. Specifically, the training processes involved two text generation tasks, generating patent titles based on abstracts and suggesting words for masked tokens in the given abstracts. The performance was better than the baseline, specifically BERT and GPT-2 on a single task. This paper indicates that multi-task learning may be a promising method for patent text generation in the future. A recent report focused on the drafting of patent sections from academic papers, such as background, summary, and description (Knappich et al., 2024). The authors highlighted the potential of LLMs in patent drafting and revealed the main challenges, including handling content repetition, retrieving contextually relevant data, and adapting the system for longer documents.
Recent studies have explored the application of LLM-based agents in various aspects of patent drafting, analysis, and management. An agent in this context refers to an autonomous system that leverages LLMs to perform specialized tasks, often in a structured, multi-agent framework where different agents collaborate to achieve complex objectives. For example, Wang et al. (2024c) proposed EvoPat, a multi-LLM-based agent designed for patent summarization and analysis. Wang et al. (2024b) introduced AutoPatent, a multi-agent framework comprising an LLM-based planner agent, writer agents, and an examiner agent, which work together for patent application drafting. Additionally, Chu et al. (2024) presented an LLM-based recommender system specifically designed to assist with patent office action responses. PatExpert (Srinivas et al., 2024) in turn suggested a meta-agent that orchestrates task-specific expert agents for various patent-related tasks, such as classification, claim generation, and summarization. It also includes a critique agent for error handling and feedback provision to enhance adaptability and reliability.
@@ 7.1 Data and Benchmarks
Additionally, benchmarks consist of labeled datasets and established metrics for performance evaluation, such as accuracy and precision. These benchmarks serve as a foundational framework that ensures researchers can assess the strengths and limitations of their methods against a standardized set of conditions to promote transparency and fairness in comparison. In the context of the patent domain, many tasks have relied on closed-source datasets, which limits the ability of other researchers to replicate and validate findings. Moreover, general text evaluation metrics may not be suitable for patent text assessment (Jiang et al., 2024c). The absence of well-defined benchmarks for these tasks renders it challenging to assess and compare the effectiveness of various models fairly and comprehensively.
@@ 7.2 Use of Large Language Models
Since texts are the key ingredient to patent documents, promising methods to analyze and generate patent texts are worth investigating. Large language models (LLMs) have dramatically changed the field and achieved remarkable performance in a wide range of tasks in the general domain (Min et al., 2023), such as information retrieval, translation, and summarization. However, few works have evaluated and investigated the usage of LLMs on patent-related tasks, leaving a large research gap. Patents include a distinct category of text, characterized by their formal language, dense technical content, and legal implications. The exploitation of LLMs to handle the intricacies of patent texts could yield considerable benefits in understanding and managing intellectual property. LLMs could identify more intricate and nuanced patterns and relationships within patent texts, facilitating the effectiveness in patent analysis tasks, such as patent classification and novelty prediction. Furthermore, LLMs could assist in generating and refining patent texts based on a comprehensive understanding of existing patent literature. For example, LLMs could assist inventors and patent attorneys in drafting initial claims, expanding descriptions, or adapting texts to align with jurisdictional requirements. The integration of LLMs into patent tasks cannot only enhance the efficiency and effectiveness of patent analysis but also improve the quality of the patent drafting process. As these models continue to evolve, LLMs appear significantly promising in the patent domain.
@@ 7.5 Multimodal Methods
Apart from text, patent documents themselves and their context include other useful information, such as drawings and citations. Multimodal models can combine multiple data types to improve the performance of patent analysis (Huang et al., 2024). Research explores model architectures such as CLIP (Radford et al., 2021) and Vision Transformers (Dosovitskiy, 2020) to bridge the gap between textual descriptions and visual data to enhance patent processing. Given the only recently increased performance gains in general multimodal methods, there are only few studies for patent applications yet, for example in patent classification (Jiang et al., 2022a) and patent image retrieval (Pustu-Iren et al., 2021). Ideally, researchers can adopt multimodal methods in other patent tasks as well. For example, multimodal models can assist in patent drafting from design figures, ensuring the textual descriptions and figures are aligned. Furthermore, multimodal methods can create more informative and intuitive visualizations of patent data, providing a clearer understanding of the patent landscape. A multimodal approach appears promising and deserves more attention.
@@ 8 Conclusion
NLP techniques play a prominent role in the automation of patent processing tasks. Text embeddings can extract word and semantic information for patent analysis. Sentence/paragraph embeddings are preferable to word embeddings because they can capture more nuanced contextual information from entire sentences or paragraphs. From the model perspective, most traditional machine-learning models, such as feature-based neural networks, support vector machines, or random forests, have already been outdated. Deep learning models, particularly convolutional and recurrent neural networks, are still used in some circumstances, for example when the computing power is limited. However, larger transformer-based models, such as the BERT and the GPT series, have almost revolutionized the field despite their excessive need for computational resources. Furthermore, scaling the model size to LLMs has led to an explosion in performance on numerous general tasks.
@@
Research has prominently focused on short text parts of patents, such as titles and abstracts, e.g., for patent analysis or generation tasks. However, these texts are the least specific texts with little information about the actual invention. Since they are highly generic and take little time during drafting, the automation of these short texts is not necessarily helpful. Among the shorter parts of patents, the claims would be an exception, defining the legal scope of patent protection. The tasks of claim generation, description generation, and LLM-based patent agents are worth investigating.
@@
The most studied and developed topics are automated patent classification and retrieval, which are routine tasks in patent offices. Transformer-based language models have shown better effectiveness in leveraging text information compared to traditional machine learning models and deep word embeddings. A more promising direction is to hybridize methods so that they can use multiple sources from patents, including texts, images, and metadata. Information extraction serves as a primer to support various patent-related applications. State-of-the-art language models can help in precise extraction to improve the performance in further applications. Furthermore, we found that the conception of novelty used in research can substantially deviate from its strict definition in patent law. Accordingly, a share of the studies may aim at predicting the patent novelty but actually generate some form of metric for oddness, potential creativity, or perceived originality. Such cases indicate the often underestimated difficulties when working with patents. Interestingly, practically no capable foundation models can deal with the particularities of the patent domain and simplify the derivation of usable tools for specific tasks. Granting prediction, litigation prediction, and patent valuation are comparably less investigated. The key acts in patent examination are almost ignored so far, including the automated formal novelty and inventiveness detection for higher quality and objectivity, deriving of arguments, and support in formulating motions.
@@
Latest language models play a role in generative tasks around patents and reflect the most recent trend towards generative artificial intelligence technology. Apart from the patent analysis tasks, the language particularities of patents also complicate generative tasks. Summarization tasks and automated translation tasks are well-established and available online. Particularly, automated translation has been studied and in production at patent offices for more than a decade. Less conspicuous yet socially beneficial tasks, such as patent simplification, receive scant attention in research. This oversight is primarily attributed to the absence of appropriate datasets. Although patent drafting, especially the sections on descriptions and claims, is an obvious task, the available research is limited and generally unsatisfactory.